Pat pays $250 to be a gym member for 2 months and $550 to be a member for 6 months. What is the monthly cost of a gym membership?
Answer by ewatrrr(24785)

(n , C)
(5, 550)
(2, 250) m = 300/3 = 100
C -250 = 100(n-2)
C = 100n + 50 (n the number of months)